You are planning for your future and want to know what the monthly payment on your...

  1. You are planning for your future and want to know what the monthly payment on your student loans will be, you have $150,000 outstanding at an interest rate of 9%. You want to pay it back over a 10 year period.
  2. You just turned 30 and want to retire in 35 years with $1,500,000. Assuming you can get a rate of return of 6%, how much money will you need to put away each month to achieve your goal?
  3. If you were to invest $50,000 from you lottery winnings today, how much will it be worth in 20 years assuming a 7% rate of return?
  4. You have been hired by a company to help then plan for a future liability, in 5 years the company needs to pay back a shareholder loan of $100,000. How much money do they need to invest today to appreciate to $100,000 in 5 years given a 4% rate of return?
  5. You have found your dream home! The sales price you agreed upon is $850,000. You have 20% to use as a down payment and will finance the rest over a 30 year period at a rate of 3.75%. How much will your monthly payments be?
  6. You are the CFO of your company, the CEO has decided that the company has outgrown its Head Quarters. The CEO wants to move in 5 years and anticipates the move will cost $1,700,000. How much money does he need to set aside today to have the money to move assuming a rate of return of 5%?
  7. You decide to stop drinking coffee and set aside the cost savings for your future, you estimate you will be able to put aside $100 per month, assuming a rate of 9% how much will you have in 10 years?

This is a Time Value of Money Assignment.
Answer #1

Solution:

a)
In the time value register put following values:
N = 10*12 = 120
FV = 0
PV = -150000
I/Y = 9/12 = 0.75
Compute the value of PMT = 1900.14
$1,900.14
b)
In the time value register put following values:
N = 5*12=60
FV = -1500000
PV = 0
I/Y = 6/12 = 0.5
Compute the value of PMT = 21499.20
$21,499.20
c)
In the time value register put following values:
N = 20
PMT = 0
PV = -50000
I/Y = 7
Compute the value of FV = 193484.22
$193,484.22
d)
In the time value register put following values:
N = 5
PMT = 0
FV = -100000
I/Y = 4
Compute the value of PV = 82192.71
$82,192.71
e)
Sales price = 850000
Down payment = 20% of sales price = .2 * 850000 = 170000
Amount left = 850000 - 170000 = 680000
In the time value register put following values:
N = 30*12=360
PV = -680000
FV = 0
I/Y = 3.75/12=0.3125
Compute the value of PMT=3149.19
$3,149.19
f)
In the time value register put following values:
N = 5
PMT = 0
FV = -1700000
I/Y = 5
Compute the value of PV = 1331994.48
$1,331,994.48
g)
In the time value register put following values:
N = 10*12=120
PMT = -100
PV = 0
I/Y = 9/12=0.75
Compute the value of FV = 19351.43
$19,351.43